# Sorting the Babble in Babel: Assessing the Performance of Language Detection Algorithms on the OpenAlex Database

### v3

@@ 2.2 Algorithms
All language detection algorithms whose performance is assessed here have been selected based on two criteria: Python implementability and completeness. The first criterion is not in itself problematic, since most language detection algorithms are implemented in Python or can be used in Python via packages specifically developed to that end. As for the second criterion, an algorithm is said to be complete if it terminates with a solution when one exists (pearl1984heuristics); in the case of the present project, any algorithm said to be complete is thus one that is able to return a linguistic prediction for each document given as input. Although this criterion does not seem too restrictive at first sight, some algorithms have a relatively low language coverage and have not been designed or trained to detect all the languages present in the OpAl database. For this and other reasons, any algorithm that was unable to return a prediction for every document submitted was excluded from subsequent analyses, regardless of its overall performance. Based on these restrictions, the different language detection algorithms selected and evaluated for this are shown and described in Table 1.
@@
CLD2 Second installment of the Compact Language Detector developed by Google for its Chrome browser, CLD2 is built on a Naïve Bayesian classifier trained on a manually constructed corpus of chosen web pages for each language (sites2025compact).
CLD3 Third installment of the Compact Language Detector designed for Chrome. It is a multi-layered neural network built on word embeddings based on character unigram, bigram, and trigram proportions and able to detect 107 languages (salcianu2025).
DetectLanguage Partly based on Google’s CLD2 Cloud translation API, this language detection web service allows its users to detect the language of a text either through a POST request or its API (detectlanguage2025).
FastText Created by Facebook’s AI Research lab for efficient learning of word representations and sentence classification, fastText’s language identification model is trained on data from Wikipedia, Tatoeba, and SETimes (joulin2016bag; joulin2016fasttext; bojanowski2016enriching). Available in both uncompressed and compressed versions, only the uncompressed and more performant version is used here.
FastSpell FastSpell complements FastText with spell-checking functionalities from Hunspell. This algorithm detects languages by first trying to determine the language of a sentence with FastText, then makes extra checks with Hunspell to determine the language more precisely (banon2024fastspell).
Langdetect A Java-based language-detection library, Langdetect is built on a Naive Bayes classifier, trained on a mix of n-gram-based language-independent character corpora extracted from Wikipedia abstracts (nakatani2010langdetect)
LangID This algorithm uses a Naive Bayes classifier trained on relative frequencies of character n-grams extracted from JRCAcquis, ClueWeb 09, Wikipedia, Reuters RCV2, and Debian i18n (lui2011cross; heafield2015language; lui2012langid)
Table 1: Description of the different language detection algorithms
For every algorithm implemented, different linguistic classification procedures were designed by using it on every metadata corpus generated, resulting in a total of 47=284*7=28 different sample partitions (117=7711*7=77 in the appendix).
@@ 2.3 Performance Assessment
To assess the predictive power of each classification technique as regards to article language, a confusion matrix was built for each of the 11 most frequent languages in the database (those identified in figure 1) by determining, for each article entry sampled, if the language detected by the algorithm and the one observed through manual annotation are in that language. In the case of English for example, this procedure consists in determining, for each annotated article, if the metadata language detected by the algorithm for the corresponding article is in English and if manual checking has determined (observed) that the article is indeed written in English.
Following this and in accordance with the procedure used in cespedes2025evaluating, true positive counts, false positive counts, and false negative counts were compiled and grouped by by classification procedure and language. Then, in order to assess the performance of each language detection procedure for each different language category, three different but complementary performance measures were implemented and used.
@@
Often used as a complement to precision, recall aims to assess how exhaustive an algorithm’s predictions are. Here, the recall score of an algorithm refers to the proportion of articles written in a given language that are identified as such by the algorithm. To take the previous example, if we were to ask an algorithm for all articles written in German, the recall score of that algorithm for German corresponds to the proportion of all articles written in German that are returned by the algorithm. Both precision and recall have true positives at the numerator, that is, the number of accurate predictions returned by the algorithm; their distinction lies at the level of the denominator, which corresponds to the total number of positive predictions (true and false positives) and relevant cases (true positives and false negatives) respectively.
@@ 2.4 Simulation
While the procedures described above provide interesting information on the performance of the various algorithms and allow them to be broken down by language and type of procedure, these descriptions can however be misleading when it comes to assessing the performance of the various procedures at the scale of the entire database, as they do not take into account robustness to uncertainty as well as the impact of data imbalance on algorithm performance evaluation. To address these crucial dimensions of algorithm performance as best as possible, a simulation based on various probabilistic models was carried out and its results presented over various configuration of precision, recall, and processing speed weights.
@@ 2.4.1 Robustness Assessment
As results presented in cespedes2025evaluating show, the quality of OpAl’s language metadata is surprisingly good; however, it it not perfect, which means that the language frequency distributions returned by the database are off to an unknown but certain extent. Also but perhaps most importantly, OpAl is an incessantly growing and changing database, as new articles are indexed and metadata of already indexed articles gets updated. The current state of the database in terms of language proportions and types of metadata is therefore not only uncertain at the present time, but also bound to change in the more or less short term. In that perspective, algorithmic robustness, that is, the ability of an algorithm to maintain ”some desired system characteristics despite fluctuations in the behavior of its component parts or its environment” (carlson2002complexity, p.2539), is bound to have a crucial impact of the quality of any linguistic classification of the database. In other words, an evaluation of linguistic classification procedures for the OpenAlex database has to take into account the ability of each procedure to that maintain good overall performance despite uncertainties in and changes to the current state of the database.
In order to best quantify database uncertainty and algorithm robustness, probabilistic confusion matrices were generated based on the sample predictions collected from the different language detection procedures. Several studies have already demonstrated the usefulness of Bayesian uncertainty modeling in the context of algorithm performance evaluation (goutte2005probabilistic; brodersen2010balanced; benavoli2017time; caelen2017bayesian; totsch2021classifier; chlaily2023measures). The probabilistic approach adopted here will consist in modeling uncertainty by treating precision, recall as well as the relative proportion of languages and metadata types in the database as random variables modeled using different conjugate distribution models. In conformity with previous probabilistic modelisations of confusion matrices and submatrices (brodersen2010balanced; totsch2021classifier), precision and recall probabilities for each algorithm, language, and metadata configuration, were modeled using posterior beta-binomial distributions with uniform priors for the confusion submatrices of each distinct database subgroup $k$; true positive counts were used as alpha parameters for both measures, while false positives and false negatives were used as beta parameters for precision and recall distributions respectively.
@@
Using these different conjugate distributional models, overall performance scores for the OpAl database were estimated through the following simulation procedure: for each classification procedure, precision and recall scores are obtained by drawing a random number from a mixture of beta-binomial distributions build on the classification outcomes for each language category, mixture whose component weights are obtained through a single draw of the language proportion model. This random sampling process is repeated 100 000100\,000 times, and scores for each measure and procedure are compiled into a single simulation dataset.
@@ 2.4.2 Data Imbalance
As is often the case in large databases, OpenAlex suffers from severe distribution skews, many of which are caused by the preponderance of English-written articles, which allegedly account for 75.175.1% of all database entries (imbalance ratio of 3:1). Such an asymmetry, apart from the question of its representativeness of the real linguistic distribution of scholarly communication on a global scale, poses a serious problem with regard to the evaluation of the linguistic classification procedures implemented for this project. As is the case with imbalanced sets in general, while the general objective is to improve recall without hurting precision, the number of false negatives for the minority classes can hardly be reduced without at the same time increasing their false positive count (he2013imbalanced). Moveover, standard classification algorithms are often biased towards the majority English class, leading to a higher misclassification rate instances belonging to minority classes, in the present case all other languages (lopez2013insight; liu2006boosting; huang2006intelligent). In the specific context of the OpenAlex database and as pointed out in cespedes2025evaluating, classifying article language in the OpenAlex database thus involves a trade-off between precision performance for English and retrieval performance for all other languages: a lower precision score for English will result in a better performance overall across all metrics, at the expense however of an underestimation of all other languages. This is all the more troublesome as the minority classes, i.e. all non-English languages in the present project, are usually those that are the most crucial from a classification perspective (lopez2013insight). In the context of the current project, it could also be argued that these minority classes are actually the most interesting ones, as they are what motivates the linguistic classification in the first place.
@@
Regarding $\beta$ values, $\beta>1$ and $\beta<1$, allows $F_{\beta,\gamma}$ to respectively confer more or less importance to recall in relation to precision, with lower and upper bounds $\beta=0.5$ and $\beta=2$ respectively conferring to recall half and twice the importance of precision. As for $\gamma$, $\gamma=0$ excludes speed from $F_{\beta,\gamma}$, $\gamma=1$ confers equal weight to both precision and speed, while lower and upper bounds $\gamma=0.5$ and $\gamma=2$ makes speed half and twice as important as precision, in that order. Given this, $\gamma=0$, allows $F_{\beta,\gamma}$ becomes formally equivalent to $F_{\beta}$, while setting $\beta=1$ and $\gamma=1$ makes $F_{\beta,\gamma}$ return the harmonic mean of precision, recall and speed. (sasaki2007truth).
Overall simulated performance for each procedure is then obtained by calculating $F_{\beta},\gamma$ scores obtained by each procedure at each simulation step for a set of various $\beta$ values ranging from $0.5$ (precision is twice as important as recall) to $2$ (recall is twice as important as precision) as well as 4 different values of gammagamma: 0 (speed is ignored altogether), 0.5 (speed is twice as important as recall), 1 (precision and speed are equally important), 2 (precision is twice as important as speed). Following these aggregative computations, two two different performance measures are then computed: maximum a posteriori estimates, which correspond to the mode of the posterior density obtained for each procedure, and win ratios, which are computed by dividing the number of times each procedure gets the best aggregated score by the total number of simulation iterations.
Finally, maximum a posteriori estimates for each procedure and every $F_{\beta,\gamma}$ weighing configuration are obtained by extracting the mode of every corresponding posterior density distribution.
@@ 3.1 Performance by Language
The various scatter plots included in Figure LABEL:fig:perf show performance scores obtained for each language and procedure. In order to simply both visualization and interpretation, only best-performing procedures in either precision or recall were considered for each algorithm. In case of equal score between two or more procedures using the same algorithm, only the best-performing procedure as regards to the other measure was kept (i.e. best recall score for procedures with equal precision score or best precision score for procedures of equal recall score). Additionally, in case of equal precision and recall scores between two or more procedures using the same algorithm, only the one based on the shorter, simpler, and faster-processable corpus was kept (with Titles ¡ Titles & Journal Names ¡ Titles & Abstracts ¡ Greedy). Finally, in cases where a procedure from a given algorithm and language obtains the best score for both measures, no other procedures for that algorithm and languages are represented. Precision and recall scores for all procedures implemented for this study can be found in appendix D.
@@
In each subplot, procedure performance is represented by points whose X and Y coordinates respectively refer to the precision and recall scores of each corresponding procedure. The different algorithms and corpus types uniquely characterizing each procedure are for their part represented by distinct colors and letter tags, in that order. Regarding the color coding of algorithms in particular, given the fact that FastText and FastSpell algorithms perform identically for most linguistic categories considered, a different and common color was used for all languages in which performance scores were identical for both algorithms (i.e. all categories but Other languages). Finally, in order to compare overall linguistic performances, harmonic averages of precision and recall scores by language for all procedures implemented for this project (those described in appendix B) are represented in each subplot by horizontal and vertical dashed lines respectively.
@@
As the different subplots show, procedure performance varies both substantially and differently across the different subplots, with each algorithm or corpus type performing more or less well depending on the specific circumstances of each language subset evaluated. Looking at procedure scores in general, language averages for precision are higher than their recall counterparts for all languages, with the sole exception of articles written in English and Other languages. Comparisons across languages also show that average precision is highest for Russian and Indonesian articles, while the lowest scores belong to articles written in English and Other languages; as for recall, mean procedure scores are highest for French, Italian, and Portuguese articles, and are lowest on average in the case of Korean, Japanese, Chinese, and Russian articles. Overall, the performances obtained by the assessed procedures in all language categories confirm the well-known inverse relationship between precision and recall: use of the two-sided Wetzels Bayes Factor shows that both performance measures are inversely correlated at a rate of $\rho=-.18$, with the observed data supporting the presence of correlation over the null hypothesis at a ratio of 127437 to 1 (Wetzels et al., 2012).
@@
With regard to the different types of corpus implemented, all corpus types have procedures with above-average precision and recall scores in at least one language. While the various filtering procedures described above prevent any in-depth analysis of performance by corpus types, much more can however be said of performances by algorithms and languages. First and foremost, the first quadrant of the plane determined by the two perpendicular lines representing the precision and recall averages include points in all cases, which means that every language includes at least one procedure which is above average in both precision and recall. Moreover, each algorithm has procedures that score above both precision and recall average in at least language: while FastText and FastSpell have at least one procedure with above-average precision and recall scores in all languages but Spanish and Russian, Spanish is the only language where the CLD3 algorithm scores above average in both precision and recall. In addition, two algorithms (Langdetect for articles in English, Chinese, and Other languages as well as CLD3 for Japanese and Other languages) have procedures whose precision and recall scores are both below language average.
@@
As regards to each measure considered separately, no single algorithm has above-average precision scores in all languages, while FastSpell and FastText are the only algorithms with at least one above-average recall score in each language category. These two algorithms are also those that stand out the most across languages, with procedures performing above precision or recall average in all languages except Spanish and Russian. In the case of French, Indonesian, and Italian articles, FastSpell and FastText are actually the only algorithm whose procedures score above average for both measures. In the case of Other language category, which is the only one in which the performance of both algorithms differ, FastText performs markedly better in terms of precision, while recall scores of FastSpell are slightly higher than those of FastText; for all other languages, the additional features of Fast Spell have no impact other than increasing the space and time complexity of the procedures involved.
@@
Finally, a note regarding English-written articles, which account for the vast majority of articles indexed in the database. FastText, FastSpell, and to a lower extent LangID are the only algorithms recording above-average scores in both precision and recall; besides these algorithms, DetectLanguage and CLD2 have procedures that score above-average in terms of recall, while CLD3 is the only other algorithm with above-average precision scores. The highest precision score is recorded by the LangID algorithm on the Greedy corpus, while the best recall performance is obtained by applying the DetectLanguage algorithm on the Titles & Journal Names corpus. However, as both procedures have relatively low scores in the case of the other measure, it can be concluded that procedures based on the FastSpell and FastText algorithms (in conjunction with the one consisting in the use of the LangID on the Titles corpus, but to a lesser extent) are those whose overall performance stand out the most for English. Given the hegemony of that language in scientific publishing and its prevalence in the database, we should also expect these procedures to perform relatively well in the large-scale simulation that follows. However, in light of the peculiarities mentioned in the Methods section regarding the precision-recall trade-offs in imbalanced datasets, the linguistic classification performance of English-written articles cannot be assessed alone, independently of those of all other languages present in the database. On the contrary, a linguistic classification of all articles indexed in the OpenAlex database can only be effective insofar at it integrates all languages and weighs its impact on each of them in a way similar in scope and aim to the simulation conducted here and whose results are presented and discussed below.

### v4

@@ 3.2 Algorithms
All LI algorithms whose performance is assessed here have been selected based on two criteria: Python implementability and completeness. In the first case, only off-the-shelf, python-based, LI systems that do not require any sort of training or parametrization by the user are considered here. The main reason for this is that since OpenAlex is Python-based, evaluating classification procedures implemented in other languages would be unjustified from an implementational perspective, as it would only raise unnecessary feasability issues. This first criterion is not problematic in any case, since most LI algorithms are implemented in Python or can be used in Python via packages specifically developed to that end. Another reason for the use of Python-based algorithms is that the OpenAlex back-end is itself Python-Powered; Large Language Models were also excluded for this reason as well for their inherent scalability limitations, which prevent their use on a large-scale database like OpenAlex. As for the second criterion, an algorithm is said to be complete if it terminates with a solution when one exists [pearl1984heuristics]; in the case of the present project, any algorithm said to be complete is thus one that is able to return a linguistic prediction for each document given as input. Although this criterion does not seem too restrictive at first sight, some algorithms have a relatively low language coverage and have not been designed or trained to detect all the languages present in the OpenAlex database. For this and other reasons, any algorithm that was unable to return a prediction for every document submitted was excluded from subsequent analyses, regardless of its overall performance.
Based on these restrictions, the different LI algorithms selected and evaluated for this are shown and described in Table 2.
@@
CLD2 Second installment of the Compact Language Detector developed by Google for its Chrome browser, CLD2 is built on a Naïve Bayesian classifier trained on a manually constructed corpus of chosen web pages for each language [sites2025compact].
CLD3 Third installment of the Compact Language Detector designed for Chrome. It is a multi-layered neural network built on word embeddings based on character unigram, bigram, and trigram proportions and able to detect 107 languages [salcianu2025].
DetectLanguage Partly based on Google’s CLD2 Cloud translation API, this LI web service allows its users to detect the language of a text either through a POST request or its API [detectlanguage2025].
FastText Created by Facebook’s AI Research lab for efficient learning of word representations and sentence classification, fastText’s language identification model is trained on data from Wikipedia, Tatoeba, and SETimes [joulin2016bag, joulin2016fasttext, bojanowski2016enriching]. Available in both uncompressed and compressed versions, only the uncompressed and more performant version is used here.
FastSpell FastSpell complements FastText with spell-checking functionalities from Hunspell. This algorithm detects languages by first trying to determine the language of a sentence with FastText, then makes extra checks with Hunspell to determine the language more precisely [banon2024fastspell].
Langdetect A Java-based language-detection library, Langdetect is built on a Naive Bayes classifier, trained on a mix of n-gram-based language-independent character corpora extracted from Wikipedia abstracts [nakatani2010langdetect]
LangID This algorithm uses a Naive Bayes classifier trained on relative frequencies of character n-grams extracted from JRCAcquis, ClueWeb 09, Wikipedia, Reuters RCV2, and Debian i18n [lui2011cross, heafield2015language, lui2012langid]
Table 2: Description of the different LI algorithms
For every algorithm assessed here, different linguistic classification procedures were designed by applying it on each metadata corpus generated, resulting in a total of 47=284*7=28 different sample partitions.
@@ 3.3 Performance Assessment
To assess the predictive power of each classification proceudre as regards to article language, a confusion matrix was built for each of the 12 different linguistic categories (the 11 most frequent languages in the database (those identified in figure 1, along with a 12th category including all other languages) by determining, for each language sampled, if the language detected by the algorithm and the one observed through manual annotation are the same. In the case of English for example, this procedure consists in determining, for each annotated article, if the metadata language detected by the algorithm for the corresponding article is in English and if manual checking has determined (observed) that the article is indeed written in English. Following this and in accordance with the procedure used in cespedes2025evaluating, true positive counts, false positive counts, and false negative counts were compiled and grouped by by classification procedure and language. Then, in order to assess the performance of each LI procedure for each different language category, three different but complementary performance measures were implemented and used.
@@
Often used as a complement to precision, recall aims to assess how exhaustive an algorithm’s predictions are. Here, the recall score of an algorithm refers to the proportion of articles written in a given language that are identified as such by the algorithm. To take the previous example, if we were to ask an algorithm for all articles written in German, the recall score of that algorithm for German corresponds to the proportion of all articles written in German that are returned by the algorithm. Both precision and recall have true positives at the numerator, that is, the number of accurate predictions returned by the algorithm; their distinction lies at the level of the denominator, which respectively corresponds to the total number of positive predictions (true and false positives) and relevant cases (true positives and false negatives).
@@ 3.4 Simulation
While the procedures described above provide interesting information on the performance of the various algorithms and allow them to be broken down by language and type of procedure, these descriptions can however be misleading when it comes to assessing the performance of the various procedures at the scale of the entire database as they do not take into account robustness to uncertainty as well as the impact of data imbalance on algorithm performance evaluation. To address these crucial dimensions of algorithm performance, a simulation based on various probabilistic models was carried out and its results presented over various configurations of precision, recall, and processing speed weighing.
@@ 3.4.1 Robustness Assessment
As results presented in cespedes2025evaluating show, the quality of OpenAlex’s language metadata is very good; however, it is not perfect, which means that the language frequency distributions returned by the database necessarily differ from the actual ones to an unknown extent. Also, OpenAlex is an incessantly growing and changing database, as new articles are indexed and metadata of already indexed articles gets updated. The current state of the database in terms of language proportions and types of metadata is therefore not only uncertain at the present time, but also bound to change in the more or less short term. In that perspective, algorithmic robustness, that is, the ability of an algorithm to maintain ”some desired system characteristics despite fluctuations in the behavior of its component parts or its environment” [carlson2002complexity, p.2539], is bound to have a crucial impact of the quality of any linguistic classification of the database. In other words, an evaluation of linguistic classification procedures for the OpenAlex database has to take into account the ability of each procedure to that maintain good overall performance despite uncertainties in and changes to the current state of the database.
In order to best quantify database uncertainty and algorithm robustness, probabilistic confusion matrices were generated based on the sample predictions collected from the different LI procedures. Several studies have already demonstrated the usefulness of Bayesian uncertainty modeling in the context of algorithm performance evaluation [goutte2005probabilistic, brodersen2010balanced, benavoli2017time, caelen2017bayesian, totsch2021classifier, chlaily2023measures]. The probabilistic approach adopted here will consist in modeling uncertainty by treating precision, recall as well as the relative proportion of languages and metadata types in the database as random variables modeled using different conjugate distribution models. In conformity with previous probabilistic modelisations of confusion matrices and submatrices [brodersen2010balanced, totsch2021classifier], precision and recall probabilities for each algorithm, language, and metadata configuration, were modeled using posterior beta-binomial distributions with uniform priors for the confusion submatrices of each distinct database subgroup $k$; true positive counts were used as alpha parameters for both measures, while false positives and false negatives were used as beta parameters for precision and recall distributions respectively.
@@
Using these different conjugate distributional models, overall performance scores for the OpenAlex database were estimated through the following simulation procedure: for each classification procedure, precision and recall scores are obtained by drawing a random number from a mixture of beta-binomial distributions build on the classification outcomes for each language category, mixture whose component weights are obtained through a single draw of the language proportion model. This random sampling process is repeated 100 000100\,000 times, and scores for each measure and procedure are compiled into a single simulation dataset.
@@ 3.4.2 Data Imbalance
As is often the case in large databases, OpenAlex suffers from severe distribution skews, many of which are caused by the preponderance of English-written articles, which allegedly account for 75.175.1% of all database entries (imbalance ratio of 3:1). Such an asymmetry, apart from the question of its representativeness of the real linguistic distribution of scholarly communication on a global scale, poses a serious problem with regard to the evaluation of the linguistic classification procedures implemented for this project. As is the case with imbalanced sets in general, the general objective is to improve recall without hurting precision; however, the number of false negatives for the minority classes can hardly be reduced without at the same time increasing their false positive count [he2013imbalanced]. Moveover, standard classification algorithms are often biased towards the majority English class, leading to a higher misclassification rate for instances belonging to minority classes, in the present case all other languages [lopez2013insight, liu2006boosting, huang2006intelligent]. In the specific context of the OpenAlex database and as pointed out in cespedes2025evaluating, classifying article language in the OpenAlex database thus involves a trade-off between precision performance for English and retrieval performance for all other languages: a lower precision score for English will result in a better performance overall across all metrics, at the expense however of an underestimation of all other languages. This is all the more troublesome as the minority classes, i.e. all non-English languages in the present project, are usually those that are the most crucial from a classification perspective [lopez2013insight]. In the context of the current project, it could also be argued that these minority classes are actually the most interesting ones, as they are what motivates the linguistic classification in the first place.
@@
Regarding $\beta$ values, $\beta>1$ and $\beta<1$ allow $F_{\beta,\gamma}$ to respectively confer more or less importance to recall in relation to precision, with lower and upper bounds $\beta=0.5$ and $\beta=2$ respectively conferring to recall half and twice the importance of precision. As for $\gamma$, $\gamma=0$ excludes speed from $F_{\beta,\gamma}$, $\gamma=1$ confers equal weight to both precision and speed, while lower and upper bounds $\gamma=0.5$ and $\gamma=2$ makes speed half and twice as important as precision, in that order. Given this, $\gamma=0$ makes $F_{\beta,\gamma}$ formally equivalent to $F_{\beta}$, $\beta=1$ and $\gamma=1$ returns the standard $F_{1}$ score, and setting $\beta=1$ and $\gamma=1$ makes $F_{\beta,\gamma}$ return the harmonic mean of precision, recall and speed. [sasaki2007truth].
Overall simulated performance for each procedure is then obtained by calculating $F_{\beta},\gamma$ scores obtained by each procedure at each simulation step for a set of various $\beta$ values ranging from $0.5$ (precision is twice as important as recall) to $2$ (recall is twice as important as precision) as well as 4 different values of gammagamma: 0 (speed is ignored altogether), 0.5 (speed is twice as important as recall), 1 (recall and speed are equally important), 2 (recall is twice as important as speed). Finally, maximum a posteriori estimates for each procedure and every $F_{\beta,\gamma}$ weighing configuration are obtained by extracting the mode of each corresponding posterior density distribution.
@@ 4.1 Performance by Language
The various scatter plots included in Figure LABEL:fig:perf show performance scores obtained for each language and procedure. In order to simply both visualization and interpretation, only best-performing procedures in either precision or recall were considered for each algorithm. In case of equal score between two or more procedures using the same algorithm, only the best-performing procedure as regards to the other measure was kept (i.e. best recall score for procedures with equal precision score or best precision score for procedures of equal recall score). Additionally, in case of equal precision and recall scores between two or more procedures using the same algorithm, only the one based on the shorter, simpler, and faster-processable corpus was kept (with Titles ¡ Titles & Journal Names ¡ Titles & Abstracts ¡ Greedy). Finally, in cases where a procedure from a given algorithm and language obtains the best score for both measures, no other procedures for that algorithm and languages are represented. Precision and recall scores for all procedures implemented for this study can be found in appendix C.
@@
In each subplot, procedure performance is represented by points whose X and Y coordinates respectively refer to the precision and recall scores of each corresponding procedure. The different algorithms and corpus types uniquely characterizing each procedure are for their part represented by distinct colors and letter tags, in that order. Regarding the color coding of algorithms in particular, given the fact that FastText and FastSpell algorithms perform identically for most linguistic categories considered, a different and common color was used for all languages in which performance scores were identical for both algorithms (i.e. all categories but Other languages). Finally, in order to compare overall linguistic performances, harmonic averages of precision and recall scores by language for all procedures implemented for this project are represented in each subplot by horizontal and vertical dashed lines respectively.
@@
As the different subplots show, procedure performance varies both substantially and differently across the different subplots, with each algorithm or corpus type performing more or less well depending on the specific circumstances of each language subset evaluated. Looking at procedure score averages for each language, precision scores are higher than their recall counterparts except for English and Other articles. Average precision is highest for Russian and Indonesian articles, while the lowest scores belong to articles written in English and Other languages; as for recall, averages are highest for French, Italian, and Portuguese articles, and lowest in the case of Korean, Japanese, Chinese, and Russian articles. Overall, the performances obtained confirm the well-known inverse relationship between precision and recall: use of the two-sided Wetzels Bayes Factor shows that both performance measures are inversely correlated at a rate of $\rho=-.18$, with the observed data supporting the presence of correlation over the null hypothesis at a ratio of 127437 to 1 (Wetzels et al., 2012).
@@
With regard to the different types of corpus implemented, all corpus types have procedures with above-average precision and recall scores in at least one language. While the various filtering procedures described above prevent any in-depth analysis of performance by corpus types, much more can however be said of performances by algorithms and languages. First and foremost, the first quadrant of the plane determined by the two perpendicular lines representing the precision and recall averages include points in all cases, which means that every language includes at least one procedure which is above average in both precision and recall. Moreover, each algorithm has procedures that score above both precision and recall average in at least one language: while FastText and FastSpell have at least one procedure with above-average precision and recall scores in all languages but Spanish and Russian, Spanish is the only language where the CLD3 algorithm scores above average in both precision and recall. In addition, two algorithms (Langdetect for articles in English, Chinese, and Other languages as well as CLD3 for Japanese and Other languages) have procedures whose precision and recall scores are both below language average.
@@
As regards to each measure considered separately, no single algorithm has above-average precision scores in all languages, while FastSpell and FastText are the only algorithms with at least one above-average recall score in each language category. These two algorithms are also those that stand out the most across languages, with procedures performing above precision or recall average in all languages except Spanish and Russian. In the case of French, Indonesian, and Italian articles, FastSpell and FastText are actually the only algorithms whose procedures score above average for both measures. Regarding the Other category, which is the only one in which the performance of both algorithms differ, FastText performs markedly better in terms of precision, while recall scores of FastSpell are slightly higher than those of FastText; for all other languages, the additional features of Fast Spell have no impact other than increasing the space and time complexity of the procedures involved.
@@
Finally, a note regarding English-written articles, which account for the vast majority of articles indexed in the database. FastText, FastSpell, and to a lower extent LangID are the only algorithms recording above-average scores in both precision and recall; besides these algorithms, DetectLanguage and CLD2 have procedures that score above-average in terms of recall, while CLD3 is the only other algorithm with above-average precision scores. The highest precision score is recorded by the LangID algorithm on the Greedy corpus, while the best recall performance is obtained by applying the DetectLanguage algorithm on Titles & Journal Names. However, as both procedures have relatively low scores in the case of the other measure, it can be concluded that procedures based on the FastSpell and FastText algorithms (in conjunction with the one consisting in the use of the LangID on the Titles corpus, but to a lesser extent) are those whose overall performance stand out the most for English. Given the hegemony of that language in scientific publishing and its prevalence in the database, we should also expect these procedures to perform relatively well in the large-scale simulation that follows. However, in light of the peculiarities mentioned in the Methods section regarding the precision-recall trade-offs in imbalanced datasets, the linguistic classification performance of English-written articles cannot be assessed alone, independently of those of all other languages present in the database. On the contrary, a linguistic classification of all articles indexed in the OpenAlex database can only be effective insofar at it integrates all languages and weighs its impact on each of them in a way similar in scope and aim to the simulation conducted here and whose results are presented and discussed below.